the resultant of two force 2P and 2P is √10P. what is angle between the force?

Grade:11

1 Answers

Aksh
6 years ago
Let the angle between the two forces be x.By vector law of addition√2P×2P+√2P×√2P+2×2P×√2Pcosx=√10P√4P×P+2P×P+4√2P×Pcosx=√10PSquaring both sides 4P×P+2P×P+4√2P×Pcosx=10P×P4√2P×Pcosx=4P×P√2cosx=1cosx=1/√2x=45°Therefore, the angle between the two forces is 45°.